OpenStudy (anonymous):

A company that manufactures smartphones makes 70% of the phones with a touch screen and 55% of the phones with HD movie capability. 30% of the phones have both a touch screen and HD movie capability. What is the probability that a smartphone chosen at random has either a touch screen or HD movie capability?

OpenStudy (anonymous):

(A) 19/20, (B) 47/50, (C) 13/20, (D) 43/50, (E) 17/20

OpenStudy (anonymous):

i would love to help but im working with probability right now and its confusing

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

That's OK thanks anyways tho!!

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

@luigi0210 would you be able to assist me?

OpenStudy (luigi0210):

Sorry, not good at this :(

OpenStudy (anonymous):

Ahhh, would you know someone who might be able too?

OpenStudy (anonymous):

Thanks anyways...

OpenStudy (anonymous):

@johnweldon1993 could you help me?

OpenStudy (johnweldon1993):

Alright reading it confused me...but let me work through it so 70% have touch screen 55% have HD capability and 30% have both "What is the probability that a smartphone chosen at random has either a touch screen or HD movie capability?" Okay....well Probability of (A or B) = P(A) + P(B) - P(A and B) so we have \[\large \frac{70}{100} + \frac{55}{100} - \frac{30}{100} = \frac{95}{100}\] Looks like both are divisible by 5...so \[\large \frac{19}{20}\]

OpenStudy (anonymous):

Thanks so much @johnweldon1993 your a lifesaver.

OpenStudy (johnweldon1993):

No problem!